Question 4 1 pts A furniture maker has an order for 88 identical tables. If the furniture maker estimates his learning curve at 78% and can reach steady production of 6 hours per table on the 3th table, how long should it take to make all the tables? Express your answer in hours to the nearest whole hour.
Answer #1

Given that the maker has a learning rate of 78%

and he takes 6 hours on 3rd table,

time taken on 1st table would be = 6/(0.782) = 9.8612 =9.9

So we have our time to make 1st table, we will find the time to complete 1st and 2nd and rest 97 table will take of 6hour each

The total time is = 9.9 + 9.9(0.78) + 9.9(0.78)2 + 9.9(0.78)3 + 97*6

= 9.9 + 7.7 + 6 + 4.7 + 582

= 610.3 hours